OpenStudy (anonymous):

A $3,500 principle earns 3% interest, compounded semiannually. After 20 years, what is the balance in the account?

OpenStudy (whpalmer4):

Use "principal", not "principle" when speaking of money. \[FV = PV(1+\frac{r}{n})^{nt}\]where \(FV\) is future value, \(PV\) is present value, \(r\) is the interest rate, expressed as a decimal, \(n\) is the number of compounding periods per year, and \(t\) is the number of years.

OpenStudy (anonymous):

I know, I've tried doing that already, but I can't seem to do the nt part at the end.

OpenStudy (anonymous):

3500(1+.03/2)^2*22 right?

OpenStudy (whpalmer4):

close, but not quite. \[nt = n*t\]

OpenStudy (anonymous):

Yeah, n = 2 because it's semiannually right?

OpenStudy (whpalmer4):

yes, but \[2*20 \ne 22\] :-)

OpenStudy (whpalmer4):

\[3500(1+\frac{0.03}{2})^{2*20}\]

OpenStudy (anonymous):

So the nt is actually 40 ?

OpenStudy (whpalmer4):

\[n = 2\]\[t=20\]\[nt = (2)(20) = 40\]

OpenStudy (anonymous):

I just put the whole thing on a calculator and got a weird number. I think I'm doing something wrong :(

OpenStudy (whpalmer4):

what was the weird number you got?

OpenStudy (anonymous):

1.03838213e-8

OpenStudy (whpalmer4):

yep, that's wrong :-)

OpenStudy (anonymous):

Lol I knew it

OpenStudy (anonymous):

Well, what I'm I doing wrong? I have all the functions and stuff :(

OpenStudy (whpalmer4):

I don't know what buttons you're pushing, or what kind of a calculator you have. I would work it like this: 0.03/2 +1 now raise that to the 40th power * 3500

OpenStudy (anonymous):

Yeah, I did it exactly like that!

OpenStudy (whpalmer4):

Well, you must have done something wrong in the process. Try it again.

OpenStudy (anonymous):

Got it!

OpenStudy (anonymous):

6349.06

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

What is the formula for depreciate?

OpenStudy (whpalmer4):

Yeah, that looks more like it. Sorry, don't recall the depreciation formula...

OpenStudy (anonymous):

Well that's okay. thanks for your help.